OpenStudy (anonymous):

if...8,14,48,100..... then what would be a next term?? (a)180 (b)176

OpenStudy (anonymous):

Ok maybe this helps 2x4 , 2x7,2x2x12,2x2x25,...

OpenStudy (anonymous):

im come to hate these questions now lol

OpenStudy (anonymous):

176 = 2x2x2x2x11 180 = 2x2x3x3x5

OpenStudy (anonymous):

well 176 suits more ...I am with 176

OpenStudy (anonymous):

i can come up with two equations <.< one that makes a work and then one that makes b work >.> they will look really really ugly though.

OpenStudy (anonymous):

please upload them ..it will definitely help

OpenStudy (anonymous):

they wont be the "correct" answer though. i havent made them yet btw, i know how though.

OpenStudy (anonymous):

its more of a "take this for not giving us enough terms to see the pattern" in your face thing lol

OpenStudy (anonymous):

lol

jhonyy9 (jhonyy9):

nice i am with 176 too

OpenStudy (anonymous):

Someone give me 5 numbers between -10 and 10

OpenStudy (anonymous):

-9 ,-3,0,2,3

myininaya (myininaya):

i was thinking 176 too just because 180 is too pretty

OpenStudy (anonymous):

hmm, switch the 0 for something else <.< i dont think i can make that work

OpenStudy (anonymous):

lol

myininaya (myininaya):

1

OpenStudy (anonymous):

lol 180 is pretty \[\large{180}\]

jhonyy9 (jhonyy9):

because 8+6=14 and 6*8=48

OpenStudy (anonymous):

k, one sec

OpenStudy (anonymous):

but how will you approach 100 with that jhonny

myininaya (myininaya):

wait!

jhonyy9 (jhonyy9):

8+14+48+100=170+6=176

myininaya (myininaya):

wait nvm

OpenStudy (anonymous):

ok see this \[3^2 -1 = 8\] \[4^2 - 2= 14\] \[7^2 -1 = 48\] \[10^2- 0 = 100\]

OpenStudy (anonymous):

the 100 makes it look bad

OpenStudy (anonymous):

ok maybe \[13^2 + 7= 169 +7 = 176\] value goes from 0 to +ve lol it sounds very bad

OpenStudy (anonymous):

\[\frac{37}{11880}(-9)^n-\frac{119}{360}(-3)^n-\frac{1}{40}(1)^n+\frac{122}{55}(2)^n+\frac{7}{8}(3)^n\] Plugging in n = 1, 2,...., 5 gives 8, 14, 48, 100, 180.

myininaya (myininaya):

lol

OpenStudy (anonymous):

keep in mind this isnt the correct answer though >.> just a trick.

myininaya (myininaya):

why can i never see a pattern anymore :(

OpenStudy (anonymous):

i know =/ i thought i was decent at spotting relationships =/

OpenStudy (anonymous):

But your relation joe it looks ....how did you derived it

OpenStudy (anonymous):

you pick as many random numbers as you need. since i wanted the value of the function to match 5 numbers i picked 5. Then you set up something like: \[c_1(r_1)^n+c_2(r_2)^n+c_3(r_3)^n+c_4(r_4)^n+c_5(r_5)^n\] each r is one of the numbers you picked. So your goal is to figure out what the constants are.

OpenStudy (anonymous):

i just create a matrix that will give me the coefficients. usually they are really ugly. But as you can see, one could create infinitely many formulas by picking different numbers.

OpenStudy (anonymous):

yeah but good one

OpenStudy (anonymous):

its what the teacher gets! giving us some sequence with no observable pattern >.>

OpenStudy (anonymous):

How can they prove my answer is wrong? they cant! i win >.>

OpenStudy (anonymous):

seriously your school is so good ... i like this thing make patterns

myininaya (myininaya):

joe is awesome i wish had his super powers

OpenStudy (anonymous):

lol, im just covering for the fact i dont know the answer to this problem :P

OpenStudy (anonymous):

Lets vote ...make wrong right(if we vote wrong more ) I am with 176

OpenStudy (anonymous):

i say 176 as well.

OpenStudy (anonymous):

ok 176 | 180 | 2 | 0

myininaya (myininaya):

well i think 180 now

jhonyy9 (jhonyy9):

176 too

myininaya (myininaya):

we couldn't find a pattern that made 176 work he can show the pattern joe found to support is answer as 180 he has no evidence for this 176

OpenStudy (anonymous):

i would like to switch my vote.

myininaya (myininaya):

lol

OpenStudy (anonymous):

176 | 180 2 2

myininaya (myininaya):

so now we can make it 50 50 so he won't be able to say he got a wrong answer from open study

OpenStudy (anonymous):

jhonny lets switch these guys are not wrong

OpenStudy (anonymous):

do we battle to the death now?

myininaya (myininaya):

we could be wrong lol who knows

myininaya (myininaya):

ishaan dont give into peer pressure

OpenStudy (anonymous):

watch the teacher come back with some linear formula >.> i'd be mad!

myininaya (myininaya):

this doesn't look linear

OpenStudy (anonymous):

yeah its not, good point.

myininaya (myininaya):

i wonder if this person is in algebra or some class close to the ground of math

OpenStudy (anonymous):

look at this

OpenStudy (anonymous):

differences 14-8 = 6 48- 14= 34 100-48=52 6,34,52

OpenStudy (anonymous):

2x3 , 17 x2 , 13x2x2

OpenStudy (anonymous):

176 - 100 = 76 19x2x2 = 76

OpenStudy (anonymous):

all prime good primes 17 , 13 ,19

OpenStudy (anonymous):

thats is not with 80

OpenStudy (anonymous):

180 -100 = 80 not possible

OpenStudy (anonymous):

it's just random

jhonyy9 (jhonyy9):

if begin with 6,8,14,48, ... ??? can we get 100 ?

myininaya (myininaya):

i give up

myininaya (myininaya):

i am walking away and not looking back

OpenStudy (anonymous):

me too, i just dont see anything =/

myininaya (myininaya):

if i don't this problem will drive me crazy

OpenStudy (anonymous):

me too ; i advise this to everyone who reads this question